# Mixed-Reality Digital Twins: Leveraging the Physical and Virtual Worlds for Hybrid Sim2Real Transition of Multi-Agent Reinforcement Learning Policies

### v5

@@ Abstract
Multi-agent reinforcement learning (MARL) for cyber-physical vehicle systems usually requires a significantly long training time due to their inherent complexity. Furthermore, deploying the trained policies in the real world demands a feature-rich environment along with multiple physical embodied agents, which may not be feasible due to monetary, physical, energy, or safety constraints. This work seeks to address these pain points by presenting a mixed-reality digital twin framework capable of: (i) selectively scaling parallelized workloads on-demand, and (ii) evaluating the trained policies across simulation-to-reality (sim2real) experiments. The viability and performance of the proposed framework are highlighted through two representative use cases, which cover cooperative as well as competitive classes of MARL problems. We study the effect of: (i) agent and environment parallelization on training time, and (ii) systematic domain randomization on zero-shot sim2real transfer across both case studies. Results indicate up to 76.3% reduction in training time with the proposed parallelization scheme and sim2real gap as low as 2.9% using the proposed deployment method.
@@ I Introduction
Connected autonomous vehicles (CAVs) are exemplars of cyber-physical systems (CPS) operating within an environment with other agents. The development and deployment of such systems present a formidable challenge due to the complexity-growth of multi-agent interactions. In such a milieu, multi-agent learning stands out as a promising avenue for developing autonomous vehicles capable of navigating complex and dynamic environments while considering the nature of interactions with their peers. Particularly, multi-agent reinforcement learning (MARL) offers the tantalizing potential of learning through self-exploration, which can potentially capture intricate cooperative/competitive multi-agent interactions.
@@
Figure 1: Mixed-reality digital twin framework for hybrid sim2real transfer of MARL systems: (1) observe, (2) decide, (3) act, (4) estimate, and (5) update.
Figure 2: Simulation parallelization: (a) depicts a snapshot of 25 cooperative MARL environments training in parallel, and (b) denotes the training time for different levels of environment parallelization. Similarly, (c) depicts a snapshot of 10$\times$2 competitive MARL agents training in parallel, and (d) denotes the training time for different levels of agent parallelization.
@@
Irrespective of the problem formulation, one of the key challenges in training MARL policies is the sample efficiency of the environment, which translates to longer training times. This is often addressed by (a) training in low-fidelity simulations that can run faster than real-time [1, 2, 3], (b) parallelizing simulations to accelerate data collection [7, 8], or (c) both [6]. Adopting the first approach usually leads to a heightened sim2real gap, as marked in the literature through simulation-only deployments or explicit remarks for real-world experiments [6]. Adopting the second approach, on the other hand, usually requires extensive computational resources [7, 8, 9], which may not be sustainable. Additionally, none of the prior works addresses the challenge of selectively isolating collision, interaction, and perception between parallelized replicas of multi-agent systems.
Another important challenge arises during the real-world deployment of MARL policies. While generic RL has studied the sim2real transfer of trained policies using techniques such as domain adaptation [10], identification [11], or augmentation [12], the sim2real transfer of MARL systems has been typically under-explored. While recent works such as [4] and [6] have tried to fill this gap, they simplified the formulation by adopting extensive observation spaces with ground truth information about the environment and employed benchmark equipment such as mocap for real-time feedback during sim2real transfer. Additionally, these works used multiple physical vehicles, albeit scaled, within a synthetically constructed physical test environment for real-world deployments, which may not always be feasible due to monetary, spatial, energy, or safety constraints.
@@
This work tries to address these two MARL pain points through the following open-source [^1] GitHub: https://github.com/autodrive-ecosystem contributions:
@@
  • Parallel Training: This work contributes a modular simulation parallelization framework, which allows selectively isolating the exteroceptive perception, collision, and interaction within or among MARL system(s).
@@
  • Sim2Real Transfer: We introduce a bi-directional digital twinning framework to immerse a limited number of physical agents within a digital environment running virtual peer(s) to evaluate the MARL policies.
@@
  • MARL Analysis: The proposed MARL case studies are benchmarked against a decentralized reactive planning algorithm to assess their efficacy. We numerically evaluate the sim2real gap to analyze the effect of systematic domain randomization introduced in this work.
The remainder of this paper is organized as follows: Section II elucidates the two MARL case studies, including their mathematical formulation. Section III describes the workflow adopted to train and deploy the MARL policies, including their sim2real transfer. Section IV analyzes the MARL training and deployment results for both case studies. Finally, Section V provides concluding remarks and future research directions.
@@ II-A2 Action Space
The Ackermann-steered vehicles were controlled using throttle and steering commands: $a_{t}^{i}=\left[\tau_{t}^{i},\delta_{t}^{i}\right]\in\mathbb{R}^{2}$. From a collision avoidance perspective, the throttle command allowed the agents to speed up or slow down, and the steering command allowed the agents to dodge their peers.
@@ II-A3 Reward Function
This function rewarded each agent with $r_{goal}=1$ for successfully traversing the intersection and penalized them proportional to their distance from the goal, represented as $k_{p}*\left\|g_{t}^{i}\right\|_{2}$, for collisions or lane-boundary violations. The agents were also continuously rewarded inversely proportional to their distance-to-goal, to negotiate the sparse-reward problem. The reward ($k_{r}$) and penalty ($k_{p}$) constants were set to $0.01$ and $0.425$ respectively.
@@ II-B1 Observation Space
Each agent collected a vectorized observation: $o_{t}^{i}=\left[v_{t}^{i},m_{t}^{i}\right]\in\mathbb{R}^{28}$. Here, $v_{t}^{i}\in\mathbb{R}^{1}$ represents the estimated forward velocity of $i$-th agent, and $m_{t}^{i}=\left[{}^{1}m_{t}^{i},^{2}m_{t}^{i},\cdots,^{27}m_{t}^{i}\right]\in\mathbb{R}^{27}$ is the LIDAR range array with 27 ranging measurements uniformly spaced 270 around each side of the heading vector, within a 10 m radius.
@@ II-B2 Action Space
The action space to control Ackermann-steered vehicles was $a_{t}^{i}=\left[\tau_{t}^{i},\delta_{t}^{i}\right]\in\mathbb{R}^{2}$. Here, the throttle command allowed the agents to optimize their speed profile, and the steering command allowed the agents to optimize their race line, overtake their peers, and avoid collisions.
@@ II-B3 Reward Function
  • GAIL Reward: The generative adversarial imitation learning (GAIL) reward [15] ${}^{g}r_{t}$ ensured that the agent optimized its actions safely and ethically by rewarding proportional to the closeness of new observation-action pairs to those from the recorded demonstrations.
@@
  • Extrinsic Reward: The objectives of lap time reduction and motion constraints were handled using an extrinsic reward function ${}^{e}r_{t}$. The agents received a reward of $r_{checkpoint}=0.01$ for passing each of the 19 checkpoints $c_{i}$ on the race track, $r_{lap}=0.1$ upon completing a lap, $r_{best\>lap}=0.7$ upon achieving a new best lap time, and a penalty of $r_{collision}=-1$ for colliding with the track bounds or peer agent (in which case both agents were penalized equally). Additionally, a continuous reward promoted higher velocities $v_{t}$.
@@ III-A Simulation Parallelization
  • Parallel Instances: Multiple simulation instances can be spun up to train families of multi-agent systems, each isolated within its own simulation instance. This is a brute-force parallelization technique, which can cause unnecessary computational overhead.
  • Parallel Environments: Isolated agents can learn the same task in parallel environments, within the same simulation instance. This method can help train single/multiple agents in different environmental conditions, with slight variations in each environment.
  • Parallel Agents: Parallel agents can learn the same task in the same environment, within the same simulation instance. The parallel agents may collide/perceive/interact with selective peers/opponents. Additionally, the parallel agents may or may not be exactly identical, thereby robustifying them against minor parametric variations.
@@ III-B Learning Architecture
We leverage the proximal policy optimization (PPO) algorithm [20] for MARL training; justification follows. PPO is an on-policy method, which is empirically equally effective as its off-policy counterparts [21]. Moreover, PPO promotes stable and efficient learning by imposing 2 complementary constraints: (a) a clipped surrogate objective to control each action probability update, and (b) a KL divergence early stopping criteria to limit overall policy change.
In terms of policy updates, cooperative MARL uses the collective experience of all agents to update a common policy. Contrarily, competitive MARL uses the independent experience of each agent to update its individual policy. Nevertheless, in both cases, the parallelized agents contribute their experiences to update their respective herd’s policy. This results in distributed sampling, which improves data collection speed and diversity, thereby increasing its correlation with the true state-action distribution and stabilizing training.
@@ III-C Domain Randomization
We leveraged the simulation parallelization architecture to introduce systematic domain randomization [12] across $k$ agent/environment replicas. This allowed us to maintain the solver consistency across simulation time steps while introducing dynamical perturbations, which is not explored in the literature. Table II hosts the detailed domain randomization parameters for the cooperative as well as competitive MARL scenarios. Particularly, since the cooperative MARL scenario was environment-parallelized, we vary the dynamics of each environment replica in this case. Contrarily, since the competitive MARL scenario was agent-parallelized, we vary the dynamics of each agent replica in this case. Additionally, in both cases, we also introduce noise in the agents’ observations and actions at each time step. Here, the parameter $\xi$ denotes the degree of domain randomization. In this work, we analyze the effect of no domain randomization (NDR), i.e. $\xi=0$, low domain randomization (LDR), i.e. $\xi=1$, and high domain randomization (HDR), i.e. $\xi=2$.
@@
III-D Hybrid Sim2Real Transfer
We propose a hybrid method for transferring the trained MARL policies from simulation to reality. The term “hybrid” specifically alludes to a mixed-reality digital twin framework, which establishes a real-time bi-directional synchronization between the physical and virtual worlds. The intention is to minimize the number of physical agent(s) and environmental element(s) while deploying and validating MARL systems in the real world. Fig. 1 (captured at 1 Hz) depicts the sim2real transfer of MARL policies discussed in this work using the proposed framework. Here, we deploy a single physical agent in an open space and connect it with its digital twin. The “ego” digital twin operates in a virtual environment with virtual peers, collects observations, and uses the trained policy to plan actions in the digital space. The planned action sequences are relayed back to the physical twin to be executed in the real world, which updates its state in reality. Finally, the ego digital twin is updated based on real-time state estimates of its physical twin to close the loop. This process is repeated recursively until the experiment is completed. This way, we can exploit the real-world characteristics of vehicle dynamics and tire-road interactions while being resource-altruistic by augmenting environmental element(s) and peer agent(s) in the digital space. This also alleviates the safety concern of the experimental vehicles colliding with each other or the environmental element(s).
@@ IV Results
We use the follow-the-gap method (FGM) [23] as a benchmark for the MARL policies discussed in this work. The choice of FGM was justified since (a) it is capable of negotiating static and dynamic obstacles, which is essential for multi-agent systems, (b) it can work with non-holonomic Ackermann-steered vehicles adopted in this work, and (c) it is a decentralized reactive algorithm making a solid case for “apples-to-apples” comparison with the MARL policies.

### v6

@@ Abstract
Multi-agent reinforcement learning (MARL) for cyber-physical vehicle systems usually requires a significantly long training time due to their inherent complexity. Furthermore, deploying the trained policies in the real world demands a feature-rich environment along with multiple physical embodied agents, which may not be feasible due to monetary, physical, energy, or safety constraints. This work seeks to address these pain points by presenting a mixed-reality (MR) digital twin (DT) framework capable of: (i) boosting training speeds by selectively scaling parallelized simulation workloads on-demand, and (ii) immersing the MARL policies across hybrid simulation-to-reality (sim2real) experiments. The viability and performance of the proposed framework are highlighted through two representative use cases, which cover cooperative as well as competitive classes of MARL problems. We study the effect of: (i) agent and environment parallelization on training time, and (ii) systematic domain randomization on zero-shot sim2real transfer, across both case studies. Results indicate up to 76.3% reduction in training time with the proposed parallelization scheme and sim2real gap as low as 2.9% using the proposed deployment method.
@@ I Introduction
Connected autonomous vehicles (CAVs) are exemplars of cyber-physical systems (CPS) operating within an environment with other agents. The development and deployment of such systems present a formidable challenge due to the increased complexity of multi-agent interactions. In such a milieu, multi-agent learning stands out as a promising avenue for developing autonomous vehicles capable of navigating complex and dynamic environments while considering the nature of interactions with their peers. Particularly, multi-agent reinforcement learning (MARL) offers the tantalizing potential of learning through self-exploration, which can potentially capture intricate cooperative/competitive multi-agent interactions.
@@
Figure 1: Proposed mixed-reality digital twin framework for hybrid sim2real transfer of MARL systems: (1) observe, (2) decide, (3) act, (4) estimate, and (5) update.
Figure 2: Simulation parallelization: (a) 25$\times$4 cooperative MARL agents analyzed for (b) computational throughput, (c) training time, and (d) sample rate across different levels of environment parallelization; (e) 10$\times$2 competitive MARL agents analyzed for (f) computational throughput, (g) training time, and (h) sample rate across different levels of agent parallelization.
@@
In either case, one of the key challenges in MARL training is longer wall-clock times, which can be accelerated [9, 10] by (a) developing sample-efficient RL algorithms [11], or (b) accelerating simulations to improve data generation rate [12]. This work focuses on the latter aspect, which in turn can be addressed by (i) training in low-fidelity simulations that can run faster than real-time [1, 2, 3], (ii) parallelizing simulations to accelerate data collection [7, 8], or (iii) both [6]. Here, adopting low-fidelity simulations usually leads to a heightened sim2real gap, as marked in the literature through simulation-only deployments or explicit remarks for real-world experiments [6]. Contrarily, adopting brute-force parallelization usually requires extensive computational resources [7, 8, 13], which may not be sustainable – this is because existing frameworks cannot “smartly” parallelize replicas of multi-agent systems by selectively isolating collision, interaction, and perception between the agents/environments.
Another challenge arises when transitioning MARL policies from simulation to reality, either for training/fine-tuning or inference/evaluation. This is due to the requirement of multiple vehicles operating in a physical test environment, which may not always be feasible due to monetary, spatial, energy, or safety constraints. Consequently, there is a need to develop novel sim2real transfer techniques for MARL based on domain adaptation [14], identification [15], or augmentation [16] methods, which have been explored in the context of single-agent RL. For example, [17] proposes an interesting approach of using pre-recorded data to approximate ego viewpoint with a synthetic vehicle rendered as an obstacle during training. However, they use a simple kinematic model for agent simulation, and require a full-scale physical ado vehicle during testing, which can be prohibitively expensive, difficult to set up, and potentially unsafe. More recently, [4, 6] have tried to address this gap, albeit partially, in the context of MARL. Particularly, [4] randomizes agent dynamics to mitigate the sim2real gap, but misses other potential domain randomization aspects like environment dynamics or agent observations/actions. Additionally, they do not consider simulation parallelization, and wait till the end of an episode to perform domain randomization. [6] proposes system identification and domain randomization to mitigate the sim2real gap. They parallelize low-fidelity simulation instances for faster training, but without considering selective isolation of perception, collisions, and interactions of the agents, and consequently, they too have to wait till the end of an episode to perform domain randomization. While these approaches are stepping stones toward MARL sim2real transfer, they both simplified the problem by adopting extensive observation spaces with ground-truth information about the environment and employing mocap for sim2real transfer, which made the experimental setup complicated and expensive, with the trained policies heavily dependent on it. Additionally, both of them (typical of many others in literature) used multiple physical vehicles, albeit scaled, in a synthetically constructed physical test environment for real-world deployments, which may not scale well.
@@
Our work tries to address these MARL pain points by proposing and offering a unified set of open-source [^1] GitHub: https://github.com/AutoDRIVE-Ecosystem/MRDT-MARL tools for multi-agent robotics problems that were previously scattered and underexplored in related fields. The key contributions of this work can be summarized as follows:
@@
  • Smart Parallelization: We contribute a modular simulation parallelization framework, which allows selectively isolating the exteroceptive perception, collision, and interaction within or among MARL system(s).
@@
  • Digital Twinning: We introduce a bi-directional mixed-reality digital twinning framework to immerse a limited number of physical agents within a digital environment running virtual peer(s) to train/evaluate MARL policies.
@@
  • MARL Analysis: The proposed MARL case studies are benchmarked against a comparable set of baseline algorithms to assess their efficacy. We numerically evaluate the sim2real gap to analyze the effect of systematic domain randomization introduced in this work.
The remainder of this paper is organized as follows: Section II elucidates the two MARL case studies, including their mathematical formulation. Section III describes the workflow adopted to train and deploy the MARL policies, including the mixed-reality digital twinning framework for sim2real transfer. Section IV analyzes the MARL training and deployment results for both case studies. Finally, Section V provides concluding remarks and future research directions.
@@ II-A2 Action Space
The Ackermann-steered vehicles were controlled using throttle and steering commands: $a_{t}^{i}=\left[\tau_{t}^{i},\delta_{t}^{i}\right]\in\mathbb{R}^{2}$. From a collision avoidance perspective, the throttle command allowed agents to speed up/down, and the steering command allowed agents to dodge their peers.
@@ II-A3 Reward Function
This function rewarded each agent with $r_{goal}=1$ for successfully traversing the intersection and penalized them proportional to their distance from the goal, represented as $k_{p}*\left\|g_{t}^{i}\right\|_{2}$, for collisions or lane-boundary violations. The agents were also continuously rewarded inversely proportional to their distance-to-goal, to negotiate the sparse-reward problem. The reward ($k_{r}=0.01$) and penalty ($k_{p}=0.425$) coefficients were tuned based on $g_{t}^{i}$ to ensure that the agents incurred a unit positive or negative reinforcement (i.e., ±$\approx\pm 1$) at the end of an average episode, to ensure stable training and avoid reward hacking.
@@ II-B1 Observation Space
Each agent collected a vectorized observation: $o_{t}^{i}=\left[v_{t}^{i},m_{t}^{i}\right]\in\mathbb{R}^{28}$. Here, $v_{t}^{i}\in\mathbb{R}^{1}$ represents the estimated forward velocity of $i$-th agent, and $m_{t}^{i}=\left[{}^{1}m_{t}^{i},^{2}m_{t}^{i},\cdots,^{27}m_{t}^{i}\right]\in\mathbb{R}^{27}$ is the LIDAR range array with 27 measurements uniformly spaced across 270 and split around each side of the heading vector, within a 10 m radius.
@@ II-B2 Action Space
The action space to control Ackermann-steered vehicles was $a_{t}^{i}=\left[\tau_{t}^{i},\delta_{t}^{i}\right]\in\mathbb{R}^{2}$. Here, the throttle command allowed the agents to optimize their speed profile, and the steering command allowed them to optimize their race line, overtake their peers, and avoid collisions.
@@ II-B3 Reward Function
  • GAIL Reward: The generative adversarial imitation learning (GAIL) reward [20] ${}^{g}r_{t}$ ensured that the agent optimized its actions safely (mimicking safe demos) and ethically (avoiding proactive reward hacking) by promoting the closeness of new observation-action pairs to those from the recorded demonstrations.
@@
  • Extrinsic Reward: The objective of lap time reduction and incorporation of motion constraints was handled using an extrinsic reward function ${}^{e}r_{t}$. The agents received a reward of $r_{checkpoint}=0.01$ for passing each of the 19 checkpoints $c_{i}$ on the race track, $r_{lap}=0.1$ upon completing a lap, $r_{best\>lap}=0.7$ upon achieving a new best lap time, and a penalty of $r_{collision}=-1$ for colliding with the track bounds or peer agent (in which case both agents were penalized equally). Additionally, a continuous reward promoted higher velocities $v_{t}$.
@@ III-A Simulation Parallelization
  • Parallel Instances: Multiple instances of the simulator application can be spun up to train families of multi-agent systems, each isolated within its own simulation instance. This is a brute-force parallelization technique, which can cause unnecessary computational overhead.
  • Parallel Environments: Isolated agents can learn the same task in parallel environments, within the same simulation instance (refer Fig. 2(a)). This method can help train single/multiple agents in different environmental conditions, with slight variations in each environment.
  • Parallel Agents: Parallel agents can learn the same task in the same environment, within the same simulation instance (refer Fig. 2(e)). The parallel agents may collide/perceive/interact with selective peers/opponents. Additionally, the parallel agents may or may not be exactly identical, thereby robustifying them against minor parametric variations.
It should be noted that parallelization beyond a certain point can hurt (i.e., point of diminishing return), wherein the increased simulation workload may slow down the training so much that parallel policy optimization can no longer accelerate it. This “saturation point” is dependent on the hardware/software configuration, and is subject to change.
@@ III-B Learning Architecture
While there is no limitation on RL algorithms to be employed within our framework, we leverage proximal policy optimization (PPO) [25] for our MARL case-studies. PPO is an on-policy method, which is empirically equally effective as its off-policy counterparts [26]. Moreover, PPO promotes stable and efficient learning by imposing 2 complementary constraints: (a) a clipped surrogate objective to control each action probability update, and (b) a KL divergence early stopping criterion to limit overall policy change.
In terms of policy updates, cooperative MARL uses the collective experience of all agents to update a common policy, a.k.a. centralized training and decentralized execution (CTDE) or multi-agent PPO (MAPPO). Contrarily, competitive MARL uses the independent experience of each agent to update its own policy, a.k.a. decentralized learning or independent PPO (IPPO), since the agents are in pure competition and not allowed to share any observations or experiences. Nevertheless, in both cases, the parallelized agents contribute their experiences to update their respective herd’s policy. This results in distributed sampling, which improves data collection speed and diversity (refer Fig. 2(d) and Fig. 2(h)), thereby increasing their correlation with the true state-action distribution, and stabilizing the training.
@@
*There are 2 identical networks: one acts as actor (policy) and the other as critic (value function).
@@ III-C Domain Randomization
We leveraged the simulation parallelization architecture to introduce systematic domain randomization [16] across $k$ agent/environment replicas. Particularly, each parallelized agent/environment was simulated with a different set of dynamical parameters during each episode (more diverse domain randomization in less time), and since the parallelized agents/environments were mutually isolated from each other, this did not interrupt numerical integration between simulation time steps (i.e., maintained solver consistency). This is a slightly different strategy compared to those explored in the literature: (i) per-interval domain randomization (e.g., [28]), which may violate solver consistency, or (ii) per-episode domain randomization (e.g., [4, 6, 29]), which may limit diversity of parameters, even if parallelized.
Table II hosts the detailed domain randomization parameters for the cooperative as well as competitive MARL scenarios. Particularly, since the cooperative MARL scenario was environment-parallelized, we vary the dynamics of each environment replica in this case. Contrarily, since the competitive MARL scenario was agent-parallelized, we vary the dynamics of each agent replica in this case. Additionally, in both cases, we also introduce noise in the agents’ observations and actions at each time step. Here, the parameter $\xi$ denotes the degree of domain randomization. In this work, we analyze the effect of no domain randomization (NDR), i.e. $\xi=0$, low domain randomization (LDR), i.e. $\xi=1$, and high domain randomization (HDR), i.e. $\xi=2$.
@@
III-D Mixed-Reality Digital Twinning
We propose a hybrid method for transferring the MARL policies from simulation to reality. The term “hybrid” specifically alludes to the mixed-reality digital twin framework, which establishes a real-time bi-directional synchronization between the physical and virtual worlds. The intention is to minimize the number of physical agent(s) and environmental element(s) while deploying and validating MARL systems in the real world. Fig. 1 (captured at 1 Hz) depicts the sim2real transfer of the trained MARL policies discussed in this work using the proposed framework while Fig. 3 (captured at 5 Hz) depicts the possibility of optionally training/fine-tuning MARL policies (e.g., if there is a significant modification in the real-world setup such as the deliberately introduced turf mat in our case) within the same framework (thereby minimizing the experimental setup while enjoying the benefits of real-world data for policy update). Here, we deploy a single physical agent in an open space and connect it with its digital twin. The “ego” digital twin operates in a virtual environment with virtual peers, collects observations, optimizes (optionally, during training/fine-tuning) and/or uses (during testing/inference) the MARL policy to plan actions in the digital space. The planned action sequences are relayed back to the physical twin to be executed in the real world, which updates its state in reality. Finally, the ego digital twin is updated based on real-time state estimates of its physical twin (estimated on board) to close the loop. This process is repeated recursively until the experiment is completed. This way, we can exploit the real-world characteristics of vehicle dynamics and tire-road interactions while being resource-altruistic by augmenting environmental element(s) and peer agent(s) in the digital space. This also alleviates the safety concern of the experimental vehicles colliding with each other or the environmental element(s), especially as operational scales and number of agents increase.
Figure 3: Immersing MARL policies in the mixed-reality digital twin framework for training, fine-tuning, or deployment.
@@ IV Results
We benchmark MARL policies discussed in this work against 3 baselines. First, we choose follow-the-gap method (FGM) [30] as a common benchmark for both cooperative and competitive tasks. Additionally, we benchmark the cooperative MARL policies against artificial potential field (APF) method [31] and timed-elastic-band (TEB) planner [32], which are common approaches for dynamic obstacle avoidance. Similarly, we also benchmark the competitive MARL policies against disparity-extender algorithm (DEA) [33] and pure behavioral cloning (PBC) [19], which are popular approaches in F1TENTH autonomous races. Finally, we also benchmark the performance of the best cooperative MARL policy before (base) and after fine-tuning (FT) in the real world to adapt to the deliberately introduced turf mat.